### v1

# Robotic Ultrasound-Guided Femoral Artery Reconstruction of Anatomically-Representative Phantoms

Abstract
Femoral artery access is essential for numerous clinical procedures, including diagnostic angiography, therapeutic catheterization, and emergency interventions. Despite its critical role, successful vascular access remains challenging due to anatomical variability, overlying adipose tissue, and the need for precise ultrasound (US) guidance. Errors in needle placement can lead to severe complications, restricting the procedure to highly skilled clinicians in controlled hospital settings. While robotic systems have shown promise in addressing these challenges through autonomous scanning and vessel reconstruction, clinical translation remains limited due to reliance on simplified phantom models that fail to capture human anatomical complexity. In this work, we present a method for autonomous robotic US scanning of bifurcated femoral arteries, and validate it on five vascular phantoms created from real patient computed tomography (CT) data. Additionally, we introduce a video-based deep learning US segmentation network tailored for vascular imaging, enabling improved 3D arterial reconstruction. The proposed network achieves a Dice score of 89.21% and an Intersection over Union of 80.54% on a newly developed vascular dataset. The quality of the reconstructed artery centerline is evaluated against ground truth CT data, demonstrating an average $L_{2}$ deviation of 0.91$\pm$0.70 mm, with an average Hausdorff distance of 4.36$\pm$1.11mm. This study is the first to validate an autonomous robotic system for US scanning of the femoral artery on a diverse set of patient-specific phantoms, introducing a more advanced framework for evaluating robotic performance in vascular imaging and intervention.
I Introduction
Femoral arterial access is a fundamental component of interventional medicine, enabling procedures such as diagnostic angiography, therapeutic catheterization, and emergent interventions. Its clinical utilization has expanded over time, with femoral access just for transcatheter aortic valve replacement in the U.S. increasing from 57.1% in 2011 to 95.3% in 2019 [1]. Despite its widespread adoption, femoral access remains one of the most challenging vascular entry sites [2], with a reported 5.9% failure rate in percutaneous aortic procedures [3]. The procedure’s complexity is compounded by anatomical variability, overlying adipose tissue, and the required accuracy in US imaging interpretation. Improper needle placement significantly increases the risk of serious complications, including retroperitoneal hemorrhage, thrombosis, and other vascular sequelae [2, 4]. Given these complexities, femoral arterial access is currently performed only by highly skilled physicians in controlled hospital environments.
Robotic systems can address aforementioned challenges by leveraging their precision and repeatability in US-guided vascular interventions. These systems can enable accurate 3D reconstruction of arterial structures through automated US image segmentation, serving either as fully autonomous platforms, or as clinician-assistive technologies. Furthermore, they present significant potential for deployment in resource-constrained environments where specialized healthcare personnel are unavailable, such as in combat zones, remote locations, or an ambulance. Although robotic US scanning and reconstruction is a mature research topic, most investigations have employed overly simplified phantom models that cannot represent the anatomical complexity encountered in clinical scenarios [5, 6], or relied on single custom-designed phantoms for both system development and validation [7, 8]. Additionally, to the best of the authors’ knowledge, no study has validated robotic femoral artery reconstruction using an independent 3D ground truth derived from a separate imaging modality than that employed by the robotic system itself [9, 10].
Figure 1: Overview of the robotic system and workflow for 3D artery reconstruction.
In this manuscript, we present an autonomous robotic system for bifurcated femoral artery scanning. The proposed system integrates US image acquisition, arterial segmentation, and 3D centerline reconstruction. In particular, this work’s main contributions are as follows: (1) demonstration of autonomous robotic scanning and 3D reconstruction of bifurcated femoral artery on five vascular phantoms created from real patient computed tomography (CT) data; (2) development of a video-based segmentation network for axial artery segmentation in US video sequences; (3) comprehensive evaluation of the reconstructed bifurcated artery against ground-truth CT data to assess reconstruction accuracy. This study is the first to validate an autonomous robotic system for US scanning of the femoral artery on a diverse set of patient-specific phantoms, introducing a more rigorous framework for evaluating robotic performance in vascular imaging and intervention.
II Related Work
II-1 Vascular Phantoms
Commercial vascular phantoms, while providing good US image quality, assume the form of geometric shapes such as rectangular prisms (CIRS Vascular Access Training Phantom model 72) or semi-cylinders (CAE Vascular Access Phantom) that oversimplify the topology of the scanning surface, and limit the transferability of the system to real human procedures. Though larger models such as the Blue Phantom’s Femoral Access model offer improved anatomical representation, they are often prohibitively expensive. Perhaps most significantly, these phantoms typically derive from a single anatomical model, limiting their representation of human variation, particularly for individuals with higher body mass indices (BMI). Recently, wearable training systems combined with US simulation software, such as the SonoSkin trainer (Simulab, Seattle, WA), garnered more attention. The systems consist of a skin-like cover that can be placed over a mannequin or person, paired with a simulated US probe that triggers predetermined US images at specific anatomical locations. While such systems enable basic examination practice, their reliance on pre-loaded images and lack of orientation-specific feedback limits their utility in developing image acquisition skills, or advancing interpretive technologies. Creating a custom vascular phantom requires developing self-healing materials and durable vessels to withstand repeated needle punctures, a technology largely protected as proprietary by commercial manufacturers. To overcome these challenges, we provide a cost-effective, simple, yet highly effective solution. We develop five novel patient-specific phantoms by modifying the CIRS model $72$ vascular phantom’s surface topology and overall orientation to better represent realistic anatomy, while maintaining the phantom’s established acoustic and mechanical properties.
II-2 Vascular Segmentation in Utrasound
Automatic segmentation of US images acquired during robotic scanning presents significant challenges due to impedance control implementation - while this strategy ensures safe human-robot interaction, it introduces variations in probe contact and pressure. The inherent limitations of US imaging, characterized by poor signal-to-noise ratio and acoustic artifacts, further complicate accurate vascular boundary delineation [11]. Earlier works in US vessel segmentation used geometric and dynamic models, which typically require seed points initialization in the initial frame [12, 13, 14]. To enhance performance, previous works incorporated additional context by using Doppler US with B-mode images [15, 16, 17]. Recently, deep learning has garnered significant attention, with most vessel segmentation studies leveraging UNet-like models that process individual US images [18, 19, 20]. With the advent of transformers for modeling long-range dependencies, several attention mechanisms have been combined with U-Net-like models to encode temporal memory from videos. However, to the best of author’s knowledge, most studies have focused on obstetrics [21], breast sonography [22], and echocardiography [23]. Comparatively limited research has addressed video-based vascular US segmentation, particularly in robotic scanning.
II-3 Robotic Ultrasound Vascular Scanning
Robotic US has emerged as a promising solution to address several drawbacks of hand-held US imaging for trauma [24, 25], abdominal [26, 27] and cardiac [28] procedures. In particular to vascular US imaging, prior research has focused on vessel segmentation, tracking, and localization. Jiang et al. [29] used segmented vascular structures to align a robotic probe perpendicular to the anatomy and later refined spatial positioning using 3D point cloud data of the scanning surface [30]. Chen et al. [31] used color Doppler with B-mode images and analyzed them through deep learning for vascular access. Merouche et al. [7] conducted the automatic tracking of artery and 3D reconstruction by initializing the cross-sectional view of the artery. Ning et al. [32] employed a weakly supervised US vessel segmentation network for probe orientation estimation, and a force-guided reinforcement learning agent to maintain optimal probe angle. However, these systems were validated on either commercial or custom-designed phantoms, which fail to fully capture the anatomical complexity of real human vasculature. While some studies have validated their approaches on animals [9] or human [32] subjects, they were primarily limited to US image acquisition with no independent ground truth for validation. In contrast, our proposed robotic US vascular scanning system is validated on patient-specific phantoms with their corresponding CT scans, a safer and representative testbed for evaluating interventional procedures.
This manuscript is structured as follows: Section III.A presents the methodology and rationale underlying our vascular phantoms development. Section III.B delineates the autonomous robotic US scanning protocol, introducing a novel video-based US segmentation network. Experimental findings from the robotic US images acquisitions are analyzed in Section IV. Finally, Section V synthesizes our conclusions and presents future research directions.
III Materials and Methods
III-A Patient-Specific Vascular Phantoms
The commercial CIRS 72 model was selected as a base vascular phantom, which contains four vessel structures: two vessels of different diameters on the sides, and a central bifurcating vessel that serves as our experimental target. To reduce detection artifacts, the blood-mimicking fluid was drained from both sides of the vessels. We utilized two CIRS phantoms: an older specimen with increased hardness due to aging, and a newer phantom with softer material properties. Five anonymized CT scans from trauma patients (four male, one female) were used as a reference for phantom surface topology creation. These were full-body CT scans, with corresponding surface meshes for the skin, skeleton, and vasculature. The patients had varying BMI, as increased adipose tissue frequently complicates femoral arterial access. Clinically, vascular access targets the common femoral artery segment between the bifurcation and inguinal ligament. Unlike commercial phantoms with parallel vessels, human femoral arteries exhibit natural curvature and angulation. To enhance anatomical fidelity, we use the Meshmixer CAD software to precisely align the phantom’s bifurcated vessel with patient-specific anatomy by matching the 3D position of the bifurcation point and the orientation of the common femoral artery (see Figure 2). While each patient provided bilateral data, symmetry was assumed, and one side per patient was randomly selected for phantom creation.
Figure 2: (a) Phantom’s vessel alignment with patient’s femoral artery achieved by matching the 3D position of the bifurcation point and the orientation of the common femoral artery. (b) Aligned phantom artery relative to patients’ femoral arteries.
Figure 3: Visualization of the conceptual framework underlying the developed phantoms, methodology employed in their fabrication, and the final patient-specific phantoms created for all five cases.
Following alignment of the phantom’s bifurcated vessel with patient femoral artery data, the phantom volume is subtracted from the patient’s body volume (Figure 3). The resulting complementary volume is used to fabricate customized 10% w/w polyvinyl alcohol (PVA) hydrogel layers that replicate patient-specific surface topology when positioned atop the commercial phantoms. Fabrication involves a three-day freeze/one-day thaw cycle before mold removal. To achieve anatomically accurate positioning representative of supine patients, support platforms are designed based on patients CT data (Figure 3) that both orient the commercial phantom correctly and ensure precise positioning of the hydrogel layer relative to the CIRS models. Figure 4 compares the designed phantom models to patient anatomies using two metrics: the shortest $L_{2}$ distance from skin surface to bifurcation point, and depth ranges for both the acceptable arterial entry zone and complete artery. While patient arteries demonstrate greater depth variation due to natural curvature compared to the phantom’s planar vessel, both exhibit comparable depth ranges and significant variability by design (12-56mm) across all cases.
Figure 4: a) Visualization of the variables used to compute depth metrics. Comparative plots showing the depth range of (b) the entire artery that can be scanned and (c) the safe access entry zone of the artery, including the bifurcation depth, for both patients and phantoms.
III-B Femoral Artery Reconstruction
Accurate 3D reconstruction of the femoral artery is essential for identifying critical anatomical landmarks, such as the bifurcation, which help estimating safe needle entry points. Accurate localization of an entry zone enhances needle insertion planning, thereby minimizing complications such as arterial trauma or suboptimal catheter positioning. This study focuses specifically on achieving high-fidelity 3D reconstruction of the femoral artery using the methodology outlined below.
III-B1 Robotic US System
The robotic system, shown in Figure 1, comprises a UR10e (Universal Robots) robotic manipulator with an integrated force/torque sensor, a convex US probe (C3 HD, Clarius Mobile Health) mounted on the robot’s end effector, and a RealSense camera (Intel Corporation) attached to a tripod.
Given the predefined anatomical placement of the femoral artery within each phantom, the system bypasses vascular localization and instead focuses on autonomous scanning from one end of the artery to the other. In a real clinical scenario, this information can be provided to the robotic system by a paramedic. In the presented pipeline, this information is extracted using the phantom’s point cloud and pre-operative CT scan. The phantom’s surface geometry is first captured using the RealSense camera, generating a point cloud representation. Using the known start and end points of the artery (from CT scan), a straight-line trajectory is computed and projected onto the phantom’s point cloud within a plane that is both perpendicular to the best-fit surface plane of the phantom and contains the trajectory line. The US probe’s orientation along the scanning path is determined by the normal vector of the best-fit plane. The resulting scanning trajectory is discretized at 1mm intervals and provided as waypoints to the robot’s controller for execution.
During scanning, the robotic system employs an impedance control strategy implemented through Universal Robots’ built-in force/torque control framework. This algorithm actively regulates the end-effector’s applied force, positional accuracy, and orientation, ensuring stable and consistent probe-surface contact. The scanning procedure was conducted with a contact force of 6 N for the older CIRS phantom and 4 N for the newer model. Throughout the scanning process, the robot’s end-effector poses and corresponding US images are recorded in a rosbag file. Synchronization is maintained using ROS2 timestamps to ensure precise temporal alignment between the positional data and acquired US frames.
To ensure seamless integration of all system components, a unified reference frame is established, as illustrated in Figure 1. The notation $T^{A}_{B}$ denotes the homogeneous transformation that maps coordinates from frame $A$ to frame $B$. The essential reference frames employed in this system include: $T^{P}_{EE}$, the transformation describing the position and orientation of the US probe relative to the robot’s end effector; $T^{C}_{RB}$, the transformation between the camera frame and the robot’s base; $T^{US}_{EE}$, the transformation between the US image and the robot’s end effector; $T^{PH}_{CT}$, the transformation aligning the physical phantom with its corresponding CAD model reconstructed from computed tomography (CT) scans.
US Image Calibration: $T^{US}_{EE}$ is determined through a point-based calibration method [33]. The procedure involves acquiring US images of a spatially fixed point in a water bath from multiple robot-configured poses, solving the constraint $B_{i}X^{-1}p_{i}=B_{j}X^{-1}p_{j}$. Calibration was performed at an US imaging depth of 100mm, maintained across all experiments. It is important to note that separate calibration matrices would be required for different imaging depths due to depth-dependent scaling variations.
Eye-to-Hand Camera Calibration: A checkerboard was rigidly mounted to the robot table, and a calibrated tool tip guided to the checkerboard’s intersection points. The tool’s position relative to the robot base is recorded while the 3D camera captures depth images, extracting corresponding point cloud data. The transformation between the camera and robot base is computed using the Least Squares method.
Phantom-to-Patient Registration: Scan start and end points are defined in patient-specific CAD models derived from the CT scans. To transfer this information to physical phantoms, an Iterative Closest Point (ICP) algorithm is used to compute the transformation between the phantom in the camera frame and its corresponding CAD model. The initial alignment is achieved by matching centroids and principal components before refining the registration with ICP.
III-B2 Vascular Segmentation NetworkVideos
We propose a deep learning network to segment vessels in US videos by extending a traditional single-frame segmentation network into a video segmentation pipeline. The network processes preceding frames using encoders to extract key and value features, which are stored in a memory module. This memory module retains the value features alongside their corresponding key features, which are subsequently accessed and integrated with query feature maps through an attention-based module. The fused features generated by this process are then passed to a decoder, which produces the final segmentation mask. An overview of the network architecture is presented in Fig. 5. The key components of the network are explained below:
Figure 5: Overview of the vascular segmentation network.
Encoder and Decoder: For the encoder and decoder components, we utilized the TransUNet architecture [34], a widely recognized network for medical image segmentation. Specifically, we selected the R50-ViT-B16 variant of TransUNet. The encoder output, initially consisting of 256 channels, was further processed through an additional encoding layer to generate key and value features with 128 and 16 channels, respectively.
Memory: The extracted key and value features are stored in a memory. During the segmentation of the query frame, temporal information must be retrieved from memory. This retrieval is accomplished by matching the key feature maps from the memory and the query frames, denoted as $\boldsymbol{{K}}_{M}^{s_{i}}$ and $\boldsymbol{{K}}_{Q}$, respectively. To optimize this process, we leverage temporal prior information as in [35], which assumes that the content in the query frame is more likely to appear in similar pixel locations within the preceding (memory) frames. This operation is implimented as correlation operation:
$\begin{split}\boldsymbol{\mathcal{X}}_{nm}^{s_{i}}(x,y)&=\boldsymbol{\mathcal{X}}(\boldsymbol{{K}}_{M}^{s_{i}},\boldsymbol{{K}}_{Q})\\ &=\boldsymbol{{K}}_{M}^{s_{i}}(x,y)^{T}\boldsymbol{{K}}_{Q}(x+n,y+m).\end{split}$ (1)
Here, $(x,y)\in[1,H]\times[1,W]$, and the pair $(n,m)\in\mathcal{R}=[-h,h]\times[-w,w]$ defines the region of search for the correlation, with $h<<H$ and $w<<W$. A softmax layer is then applied to convert the correlation maps into probability maps through
$\boldsymbol{\mathcal{P}}_{nm}^{s_{i}}(x,y)=\frac{\exp(\boldsymbol{\mathcal{X}}_{nm}^{s_{i}}(x,y))}{\sum_{s_{i}\in\mathcal{S}}\sum_{(n,m)\in\mathcal{R}}\exp(\boldsymbol{\mathcal{X}}_{nm}^{s_{i}}(x,y))}.$ (2)
The probability maps represent the attention weights, which are used to weigh both the spatial and temporal information stored in the memory. The final output tensor from the memory can be expressed as
$\begin{split}\tilde{\boldsymbol{{V}}}_{M}(x,y)=\sum_{s_{i}\in\mathcal{S}}\sum_{(n,m)\in\mathcal{R}}\boldsymbol{\mathcal{P}}_{nm}^{s_{i}}(x,y)\boldsymbol{{V}}_{M}^{s_{i}}(x+n,y+m),\end{split}$ (3)
which represents the weighted sum of all memory values $\boldsymbol{V}_{M}^{s_{i}}$ across both the temporal domain and spatial domain.
Fusion: The fusion module combines the $\boldsymbol{V}_{Q}$ and $\tilde{\boldsymbol{{V}}}_{M}$ to construct a relevant input for the decoder to carry out the segmentation of the query frame. To achieve this, an attention branch is assigned to each fusion input, as in [35]. These attention branches, equipped with sigmoid-activated gates, regulate the flow of information by controlling which features - either from the query or from memory - are passed to the decoder. The fusion module can be mathematically expressed as
$\begin{split}\boldsymbol{F}(\boldsymbol{{V}}_{Q},\tilde{{\boldsymbol{V}}}_{M})&=\sigma(\boldsymbol{A}_{Q}(\boldsymbol{{V}}_{Q},\boldsymbol{{V}}_{M}))\otimes\boldsymbol{E}_{Q}(\boldsymbol{{V}})\\ &+\sigma(\boldsymbol{A}_{M}(\boldsymbol{{V}}_{Q},\boldsymbol{{V}}_{M}))\otimes\boldsymbol{E}_{M}(\boldsymbol{{V}}),\end{split}$ (4)
where $(\boldsymbol{E}_{Q},\boldsymbol{A}_{Q})$ and $(\boldsymbol{E}_{M},\boldsymbol{A}_{M})$ are the learned convolution layers of the Query attention and Memory attention, respectively. The dimension of the fusion module output is identical to that of the encoder output.
Dataset: US images were acquired using the Clarius C3 HD scanner, with a hydrogel block placed over only the old CIRS phantom. This block matches the CIRS phantom in length and width but a varying depth. This is done to keep the physics of ultrasound image same as representative dataset from developed phantoms. Scans were performed along all four vessels at a 100mm depth, capturing a total of 4,802 images across 10 continuous scan sequences. Vessel structures were manually segmented by two annotators using the VIA tool [36]. The dataset was partitioned into training, validation, and testing sets in an 80:10:10 ratio (3,838, 476, and 488 images, respectively), ensuring proportional sequence representation across all subsets. Additional dataset statistics are provided in Table I.
TABLE I: Overview of the acquired vascular US dataset on Phantom 1.
Sequence ID 1 2 3 4 5 6 7 8 9 10
US Frames $313$ $453$ $428$ $428$ $343$ $594$ $560$ $554$ $662$ $467$
III-B3 Vascular Centerline Estimation
Following robotic US image acquisition and vessel segmentation, postprocessing is implemented to mitigate noise and artifacts. False positive detections are eliminated by excluding centroids located beyond ±85 pixels from the image center, corresponding to a 45mm width threshold (based on a 0.2694 mm/pixel scale factor at 100mm scanning depth). This threshold is empirically derived through analysis of femoral artery characteristics across the five patient models, where the maximum observed femoral artery bifurcation width was 39.68mm bilaterally. To account for variability, a 12.5% margin is incorporated, adjusting the threshold from 40 to 45mm. Contour detection is applied to the refined segmentation mask, retaining only nonzero-area contours. An ellipse is fitted to each valid contour to approximate vessel geometry, recording its centroid and area. Centroids are then transformed from pixel to 3D world coordinates using the US calibration matrix, with their stacked positions forming the reconstructed artery centerline. CT scans of both CIRS phantoms are acquired, and two distinct CAD models of the bifurcated vessels are generated. The vessels’ centerline is then extracted using 3D Slicer and the Vascular Modeling Toolkit [37], discretized at 1mm intervals to serve as the ground truth.
III-B4 Evaluation Metrics
To evaluate reconstruction accuracy, the detected centerline is aligned to the corresponding ground truth using ICP. Spatial accuracy is quantified by computing the Hausdorff distance and $L_{2}$ norm between the reconstructed centroids and the ground truth centerline. The Hausdorff distance $H(A,B)$ between two point sets $A$ and $B$ is defined as:
$H(A,B)=\max\{\sup_{a\in A}\inf_{b\in B}||a-b||,\sup_{a\in A}\inf_{b\in B}||b-a||\},$ (5)
where |||||$\cdot$||| | ⋅ | | is the Euclidean distance. The Hausdorff distance in this case is computed from the reconstructed phantom centroids to the discretized ground truth centerline, since the reconstructed centerline is more sparse compared to the ground truth. The $L_{2}$ norm is computed as:
$L_{2}=\sqrt{\sum^{N}_{i=1}||p_{i}-q_{i}||^{2}},$ (6)
where $p_{i}$ and $q_{i}$ represent the closest corresponding points on the reconstructed and ground truth centerlines, respectively, and $N$ is the total number of points.
Results are reported for both postprocessed detected centroids, and manually filtered centroids that exclude points not corresponding to the vessel. This evaluation does not assess the quality of US image segmentation, which is evaluated separately, but rather focuses on quantifying potential distortions in the 3D reconstruction of the artery centerline using our approach
Figure 6: (a) Qualitative results of the network’s segmentation on the vascular test set. Arterial 3D centerline reconstructions (red) overlaid on centerlines from the five phantoms (blue) using the older CIRS model (b) after and (c) before data filtering.
IV Results and Discussion
IV-A Vascular Segmentation in US Images
The network was trained on an NVIDIA GeForce RTX 4090 GPU (24GB) using Python $3.8$, PyTorch $2.4$, and CUDA $12.1$. The Adam optimizer was initialized with a learning rate of $2\times 10^{-4}$, adjusted by a factor of $0.8$ upon plateau detection, with a minimum threshold of $1\times 10^{-9}$. A batch size of $16$ was used. The input consisted of video sequences incorporating the previous $3$ frames ($s_{f}=3$). The TransUNet model was initially trained from scratch, with pretrained encoder weights subsequently transferred to the video-based model. A correlation window size of $5$ and a memory sequence length of $4$ were used for feature fusion. US images and corresponding ground truth labels were resized to $256\times 256$ pixels using bilinear and nearest-neighbor interpolation, respectively. Data augmentation techniques included random horizontal flipping, contrast adjustment via color jitter, rotations within ±$\pm 30^{\circ}$, and gamma correction with values randomly sampled between $1.0$ and $2.5$. The network was optimized using a joint Dice and Binary Cross-Entropy loss function for a maximum of $50$ epochs.
The network’s segmentation performance was assessed using the Dice Coefficient (DC) and Intersection over Union (IoU). On the training set, the model achieved a DC of 91.28% and an IoU of 83.95%, while on the test set, it attained a DC of 89.21% and an IoU of 80.54%, demonstrating good segmentation accuracy and generalization. The minimal performance disparity between training and test sets suggests that the model effectively mitigates overfitting. Figure 6(a) presents qualitative results from the test set, illustrating predicted segmentation masks overlaid on input images. The close correspondence between predicted and ground-truth masks further validates the model’s robustness in arterial segmentation. Due to the absence of standardized benchmark datasets for vascular segmentation, direct comparative analysis with existing methodologies presents inherent challenges. Nevertheless, for contextual reference, Morales et al. [38] reported an IoU of 77.9% on a medical phantom using their proposed technique, and an average IoU of 63.8% on pig US data.
IV-B Vascular Centerline Reconstruction
Examples of the reconstructed centerline overlaid on the ground truth centerline for the bifurcated arteries of the first (older) phantom are shown in Figure 6(b)-(c). The figure includes results both before and after manually removing outliers corresponding to erroneously detected center points by the network. Additionally, violin plots illustrating the $L_{2}$ errors are shown in Figure 7, with corresponding metrics such as the mean, standard deviation, and Hausdorff distance reported in Table II.
Figure 7: Violin plots of $L_{2}$-error in centerline reconstruction of patients (P1 through P5) for different CIRS phantoms (a) PH1 and (b) PH2.
TABLE II: Table presenting the mean error, standard deviation, and Hausdorff metrics for the different CIRS phantoms (PH1 and PH2) and patients (P1 through P5) with and without removal of erroneous datapoints, and the corresponding average values per CIRS phantom.
Mean $L_{2}$ Err. (mm) Std (mm) Hausdorff (mm)
PH 1 P1 Unfiltered 2.957 6.410 41.77
Filtered 0.741 0.463 3.046
P2 Unfiltered 3.503 7.613 58.67
Filtered 1.105 0.788 6.059
P3 Unfiltered 3.819 7.026 31.96
Filtered 0.890 0.815 5.241
P4 Unfiltered 1.711 4.777 49.56
Filtered 0.955 0.821 3.820
P5 Unfiltered 1.223 2.449 21.71
Filtered 0.887 0.629 3.647
Avg Unfiltered 2.642 5.655 40.73
Filtered 0.915 0.703 4.362
PH 2 P1 Unfiltered 2.471 4.950 15.10
Filtered 0.746 0.836 6.216
P2 Unfiltered 4.933 7.923 19.28
Filtered 1.629 1.038 6.355
P3 Unfiltered 2.358 4.349 37.68
Filtered 1.343 1.183 6.922
P4 Unfiltered 6.587 6.297 34.71
Filtered 1.645 1.035 4.743
P5 Unfiltered 1.533 2.907 22.72
Filtered 1.151 1.249 5.499
Avg Unfiltered 3.576 5.285 25.89
Filtered 1.302 1.068 5.947
The 3D points reconstructed from the segmented axial US images show strong spatial alignment with the ground truth centerlines extracted from CT scans. Similar consistency was observed in results obtained from the second CIRS model, demonstrating the robustness of our proposed method. These findings indicate that the transformation from the segmented US space to the 3D world frame preserves spatial accuracy without introducing distortions - a critical requirement for reconstructing small-scale vascular structures such as the femoral artery, which can have diameters as small as 4 mm.
The importance of filtering out erroneous centroids is best illustrated from Figure 6. The unfiltered data shows a naturally wider spread compared to that of the filtered data, making it challenging to adequately assess the reconstruction quality. The Hausdorff errors are understandably larger than the $L_{2}$ errors, as the Hausdorff metric in this context reports the largest $L_{2}$ error within the dataset (as reflected by the maximum values in the violin plots). The average filtered $L_{2}$ and Hausdorff errors for the first vascular phantom (0.915 and 4.362mm, respectively) are smaller than those of the second phantom (1.302 and 5.947mm), as shown in Table II. This discrepancy can be attributed to the poorer quality of the US images for the second phantom, leading to false positives in the segmentation. It is important to highlight that the network was trained exclusively on the first CIRS model. While both phantoms are theoretically identical, their appearance under US differs in practice due to material variations and imaging artifacts. Additionally, the smaller bifurcated vessel in the second phantom, with its reduced diameter, was frequently missed by the network for both phantoms, as it became more difficult to visualize it, especially with phantoms of patients with higher BMI. This resulted in a more sparse detection of centroids for that vessel. The issue may have been further exacerbated by the use of a convex US probe. While linear probes are typically preferred for vascular scanning, a convex probe was used due to availability constraints. As a result, probe contact with the phantom did not always ensure full transducer-surface coupling, leading to suboptimal imaging. This effect was particularly pronounced in scenarios where the natural body contours would necessitate different probe orientations, even for a skilled clinician.
Accurate US segmentation alone does not guarantee realistic 3D reconstruction; however, our approach achieves high-fidelity reconstructions even in complex phantoms with irregular topology, including arteries located as deep as 69mm in patients with higher BMI.
V Conclusions and Future Work
While robotic US for vascular access has seen significant advancements, prior studies have not validated vessel 3D reconstruction on anatomically representative phantoms. In this work, we present a robotic US-guided system for femoral artery reconstruction on a diverse set of 5 patient-specific phantoms. Our approach achieves high-fidelity vessel reconstructions even in complex phantoms with irregular topology, including deeper arteries in patients with higher BMI. Additionally, we adapted a video-based segmentation network for axial artery segmentation in US video sequences. The reconstructed artery centerline exhibited strong agreement with patient-specific CT data, the clinical gold standard for vascular imaging. These results represent a step forward in enhancing robotic US-assisted vascular access procedures, improving their accuracy and potential for clinical translation.
Future work will focus on extending this framework to evaluate the robotic system’s capability for fully autonomous vascular needle insertion, including incorporating autonomous identification of the initial and final scanning points, and identification of a safe needle entry point from the reconstructed bifurcated artery. Comprehensive validation studies will also be conducted, including experiments on animal models and fresh cadaver specimens, to rigorously evaluate the system’s performance in realistic clinical environments.